OpenStudy (anonymous):

A line contains the points (8, 9) and (–12, –7). Using point-slope form, write the equation of the line that is parallel to the given line and that passes through (–5, –15).

OpenStudy (ilovepie):

Alright, first, we know that we want an equation of a line that's parallel to another. In order to be parallel, the both lines MUST have the same slope

OpenStudy (ilovepie):

The first step here would be to find the slope of a line going through the points (8,9) and (-12,-7). Do you know how to do that or do you need help?

OpenStudy (anonymous):

Help please

OpenStudy (ilovepie):

The formula for slope is m=y2-y1/x2-x1. Which is the change in y over change in x. Just plug everything in, and you should get a slope of -16/-20 which can be simplified to 1/8

OpenStudy (ilovepie):

sorry, slope of 4/5

OpenStudy (ilovepie):

Since you want a parallel equation, both must have the same slope. The equation for point slope form is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (ilovepie):

Can you continue solving or do you need help?

OpenStudy (anonymous):

help please

OpenStudy (ilovepie):

The final equation would look like this: \[y+16=\frac{ 4 }{ 5}(x+5)\]

OpenStudy (ilovepie):

Sorry, typo. Instead of +16, the answer is +15

OpenStudy (anonymous):

Is that it??

OpenStudy (ilovepie):

yes, that's the final equation. But remember that i made a typo while writing the equation, so change the 16 to a 15.

OpenStudy (anonymous):

Oh okay thank you soo much